# On planar sections of the dodecahedron

### v2

@@ Abstract.
In the analysis of three-dimensional biological microstructures such as organoids, microscopy frequently yields two-dimensional optical sections without access to their orientation. Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections. For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq V$, and ask whether this statistic determines $X$ up to isometry. We show that this is not the case: there exist two non-isometric $7$-element subsets with identical planar statistics.
As a consequence, there exists two polytopes in $\mathbb{R}^{3}$, whose distribution of isometry classes of two-dimensional intersections is identical, while the polytopes are not themselves isometric. This result is an analogue of classical non-uniqueness phenomena in geometric tomography.
@@ 1. Introduction
In the second step we again encode this discrete data into convex bodies by a local modification along selected faces. Starting from $D$, we cut off a small congruent triangular pyramid at each vertex $v\in X$, obtaining a polytope $K_{X}$ for $X\subseteq V$. When an affine plane $H$ meets $D$ near some vertices, the local shape of the section $K_{X}\cap H$ in a neighbourhood of those vertices is completely determined, up to isometry, by the configuration $\Pi\cap X$, where $\Pi$ is the nearby vertex–plane through the intersected vertices. The distribution of isometry classes of planar sections of $K_{X}$ can therefore be written as a convex combination of contributions coming from the finitely many planar intersection types $\Pi\cap X\subset V$, with coefficients prescribed precisely by the planar statistic $\mathrm{PS}(X)$.
@@ 2. Main results and model
For $X\subset V$ the planar statistic $\mathrm{PS}(X)$ is the multiset, over all planes $\Pi$ with $|\Pi\cap V|\geq 3$, of planar congruence classes of subset inclusions $\Pi\cap X\subseteq V$, seen as a point sets in $\mathbb{R}^{3}$. More precisely, let $D\subset\mathbb{R}^{3}$ be the dodecahedron as above with vertex set $V$. A vertex-plane is an affine plane $\Pi\subset\mathbb{R}^{3}$ with $|\Pi\cap V|\geq 3$, and we denote by
@@
For a subset $X\subset V$ and a vertex-plane $\Pi\in\mathcal{H}$ we consider the inclusion of finite subsets $\Pi\cap X\subseteq V$. Two vertex-planes $\Pi,\Pi^{\prime}\in\mathcal{H}$ are called planarly congruent with respect to $X$ if there exists a Euclidean isometry $g\in\mathrm{Isom}(\mathbb{R}^{3})$ such that $g(V)=V$ and $g(\Pi\cap X)=\Pi^{\prime}\cap X$. We denote the corresponding equivalence relation on $\mathcal{H}$ by $\Pi\sim_{X}\Pi^{\prime}$ and write $\Pi]_{X}$ for the equivalence class of $\Pi$.
@@
The planar statistic of $X\subset V$ is the multiset $\mathrm{PS}(X)$ of planar congruence classes $\Pi]_{X}$ as $\Pi$ ranges over all vertex-planes in $\mathcal{H}$, counted with multiplicity. In other words, $\mathrm{PS}(X)$ records, for each vertex-plane $\Pi$, the isometry type in $\mathbb{R}^{3}$ of the inclusion $\Pi\cap X\subseteq V$, and forgets the actual position of $\Pi$.
@@
It is a basic observation that the planar statistic of an image of $X$ under an isometry of $V$, orientation preserving or not, is the same as the one for $X$; i.e. isometric subsets of $V$ yield the same planar statistics. In view of the applications in shape recognition described in the introduction, our initial hope was that the planar statistics $\mathrm{PS}(X)$ allows to recover $X$, but we show that this is not the case. Our main result is the following theorem:
@@ 3. Convex bodies
We will now refine the construction and show that $S,T$ can be used to construct three-dimensional polytopes $K_{S}$ and $K_{T}$ that are not isometric, but their distribution of isometry classes of planar intersections is identitical. Let us first explain more precisely, what that means: We denote by $\mathcal{P}$ the space of compact planar convex sets modulo Euclidean isometries. For each closed convex set $K\subset\mathbb{R}^{3}$, we consider the natural map
@@
where $\mathcal{A}(3,2)$ is the affine Grassmannian of planes in $\mathbb{R}^{3}$. Note that $\mathcal{A}(3,2)$ carries a standard infinite rotation- and translation-invariant measure $\lambda$ and we call the Borel measure $(\Phi_{K})_{*}(\lambda)$ on $\mathcal{P}$ the distribution of isometry classes of planar sections. Our normalization of $\lambda$ is in such a way that $(\Phi_{K})_{*}(\lambda)(\mathcal{P})={\rm vol}(K)$ for all compact convex bodies.
@@
There exists three-dimensional polytopes $K,L\subseteq\mathbb{R}^{3}$, such that

### v3

@@ Abstract.
In the analysis of three-dimensional biological microstructures such as organoids, microscopy frequently yields two-dimensional optical sections without access to their orientation. Motivated by the question of whether such random planar sections determine the underlying three-dimensional structure, we investigate a discrete analogue in which the ambient structure is the vertex set of a Platonic solid and the observed data are congruence classes of planar intersections. For the regular dodecahedron with vertex set $V$, we define the planar statistic of a subset $X\subseteq V$ of vertices as the distribution of isometry types of inclusions $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and ask whether this statistic determines $X$ up to isometry. We show that this is not the case: there exist two non-isometric $7$-element subsets with identical planar statistics.
As a consequence, there exist two polytopes in $\mathbb{R}^{3}$, whose distribution of isometry classes of two-dimensional intersections is identical, while the polytopes are not themselves isometric. This result is an analogue of classical non-uniqueness phenomena in geometric tomography.
@@ 1. Introduction
In the second step we again encode this discrete data into convex bodies by a local modification at selected vertices. Starting from $D$, we cut off a small congruent triangular pyramid at each vertex $v\in X$, obtaining a polytope $K_{X}$ for $X\subseteq V$. When an affine plane $H$ meets $D$ near some vertices, the local shape of the section $K_{X}\cap H$ in a neighbourhood of those vertices is completely determined, up to isometry, by the configuration $\Pi\cap X$, where $\Pi$ is the nearby vertex–plane through the intersected vertices. The distribution of isometry classes of planar sections of $K_{X}$ can therefore be written as a convex combination of contributions coming from the finitely many planar intersection types $\Pi\cap X\subset\Pi\cap V\subseteq V$, with coefficients prescribed precisely by the planar statistic $\mathrm{PS}(X)$.
@@ 2. Main results and model
For $X\subset V$ the planar statistic $\mathrm{PS}(X)$ is the multiset, over all planes $\Pi$ with $|\Pi\cap V|\geq 3$, of planar congruence classes of subset inclusions $\Pi\cap X\subseteq V$, seen as point sets in $\mathbb{R}^{3}$. More precisely, let $D\subset\mathbb{R}^{3}$ be the dodecahedron as above with vertex set $V$. A vertex-plane is an affine plane $\Pi\subset\mathbb{R}^{3}$ with $|\Pi\cap V|\geq 3$, and we denote by
@@
For a subset $X\subset V$ and a vertex-plane $\Pi\in\mathcal{H}$ we consider the inclusion of finite subsets $\Pi\cap X\subseteq V$. Two vertex-planes $\Pi,\Pi^{\prime}\in\mathcal{H}$ are called planarly congruent with respect to $X$ if there exists a Euclidean isometry $g\in\mathrm{Isom}(\mathbb{R}^{3})$ such that $g(V)=V$, $g(\pi\cap V)=\Pi^{\prime}\cap V$ and $g(\Pi\cap X)=\Pi^{\prime}\cap X$. We denote the corresponding equivalence relation on $\mathcal{H}$ by $\Pi\sim_{X}\Pi^{\prime}$ and write $\Pi]_{X}$ for the equivalence class of $\Pi$.
@@
The planar statistic of $X\subset V$ is the multiset $\mathrm{PS}(X)$ of planar congruence classes $\Pi]_{X}$ as $\Pi$ ranges over all vertex-planes in $\mathcal{H}$, counted with multiplicity.
In other words, $\mathrm{PS}(X)$ records, for each vertex-plane $\Pi$, the isometry type in $\mathbb{R}^{3}$ of the inclusion $\Pi\cap X\subseteq\Pi\cap V\subseteq V$, and forgets the actual position of $\Pi$.
@@
It is a basic observation that the planar statistic of an image of $X$ under an isometry of $V$, orientation preserving or not, is the same as the one for $X$; i.e. isometric subsets of $V$ yield the same planar statistics. In view of the applications in shape recognition described in the introduction, our initial hope was that the planar statistics $\mathrm{PS}(X)$ allows one to recover $X$, but we show that this is not the case. Our main result is the following theorem:
@@ 3. Convex bodies
We will now refine the argument and show that $S,T$ can be used to construct three-dimensional polytopes $K_{S}$ and $K_{T}$ that are not isometric, but their distribution of isometry classes of planar intersections is identitical. Let us first explain more precisely, what that means: We denote by $\mathcal{P}$ the space of compact planar convex sets modulo Euclidean isometries. For each closed convex set $K\subset\mathbb{R}^{3}$, we consider the natural map
@@
where $\mathcal{A}(3,2)$ is the affine Grassmannian of planes in $\mathbb{R}^{3}$. Note that $\mathcal{A}(3,2)$ carries a standard infinite rotation- and translation-invariant measure $\lambda$ and we call the Borel measure $(\Phi_{K})_{*}(\lambda)$ on $\mathcal{P}$ the distribution of isometry classes of planar sections. Our normalization of $\lambda$ is in such a way that $(\Phi_{K})_{*}(\lambda)(\mathcal{P}\setminus\{\varnothing\})=W_{1}(K)$ for a compact convex set $K$, where $W_{1}$ is the first quermassintegral.
@@
There exist three-dimensional polytopes $K,L\subseteq\mathbb{R}^{3}$, such that